Find all points on the y-axis that are 6 units from the point (4, -3)
Answer by stanbon(75887)

Find all points on the y-axis that are 6 units from the point (4, -3)
--------
Equation of circle with center at (4,-3) and radius = 6.
----
(x-4)^2 + (y+3)^2 = 36
----
To find points on the y-axis, let x = 0.
------
(-4)^2 + (y+3)^2 = 36
------
(y+3)^2 = 20
----
y+3 = sqrt(20) or y+3 = -sqrt(20)
---
y = -3+2sqrt(5) or y = -3-2sqrt(5)
----------
Points: (0, -3+2sqrt(5)) and (0,-3-2sqrt(5))
==============================
Cheers,
Stan H.